Evaluate the integral 1+2 2017 de Note: Use an upper-case C for the constant of integration.Let f(2) be a quadratic function such that f(0) = 6 and f(2) dac is a rational function. Determine the value of f(o). f(0) =
Answer #1

2 Ita da Put I- x 2 2 Coro 2 sinro x = cosao dx=-2sin 20 do 1+2= 1+ COS 20 cosro 1-2 = 1 - CR20 asino (-esinzo) do 2 CBO žsin